OpenStudy (anonymous):

What are the solutions of the equation 2x² = 8? Use a graph of the related function what are the coordinates of the parabola??

OpenStudy (anonymous):

Use geogebra to graph the equation after rearanging it like this \[y=2x ^{2}-8\] http://www.geogebra.org/cms/ Solve the equation using the quadratic equation. \[(-b \pm \sqrt{b ^{2}-4ac})/2a\] In your equation. a=2 b=0 c=-8 Which becomes \[-0/(2)(2)\pm (\sqrt{(0)^{2}-(4)(2)(-8)})/(2)(2)\]Which now becomes\[\pm(\sqrt{64})/4\] So \[x=\pm2\] check \[y=(2)(-2)^{2}-8\]\[y=0\] So the solution checks for x=-2 Repeat for x=2